Please help me with my math word problem. I know that I need to have to sets of problems to make it a system of equations, but I do not know how to set it up. Here is the problem:
A man rows a boat 300 feet upstream against a constant current in 10 minutes. He then rows 350 feet downstream (with the same current) in 7 minutes.

Answer by mananth(16949)

A man rows a boat 300 feet upstream against a constant current in 10 minutes. He then rows 350 feet downstream (with the same current) in 7 minutes.
..
The problem is incomplete
...
let speed of boat in still water be x feet/minute
speed of current be y feet /minute
upstream speed = x-y
down stream speed = x+y
..
distance upstream = 300 feet
distance downstream = 350 feet.
...
time dowm stream = time upstream
..
300/(x-y) =10
10x-10y 300
/10
x-y = 30.................1
350/(x+y)=7
350 = 7x+7y
/7
50 = x+y..................2
add 1 & 2
80 = 2x
/2
x= 40 feet / minute .... speed of boat in still water
y =10 feet/minute ..... speed of current.
..
m.ananth@hotmail.ca